$SPAD/src/input op.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool op.output
)set message test on
)set message auto off
)clear all
--S 1 of 2
abs(x)
--R
--R
--R   (1)  abs(x)
--R                                                    Type: Expression(Integer)
--E 1
--S 2 of 2
eval(%,x=-3.4)
--R
--R
--R   (2)  3.4
--R                                                      Type: Expression(Float)
--E 2
)spool
)lisp (bye)

———-

References^{[1] nothing}

4